### v1

@@ 3 Discussion
This study also has some limitations. While we established a most comprehensive benchmark covering 39 tasks and demonstrated that the GPFM was the most generalizable foundation model, the external validation was limited to only seven tasks. This set of external validation tasks could be expanded in future work to further assess the generalizability of the models. It is worth highlighting that, across the seven external validation tasks, the GPFM achieved an average rank of 1.57, while the second-best performing model, UNI, achieved an average rank of 2.79. This result is consistent with the conclusions drawn from the experiments across the entire benchmark, highlighting the superior generalizability of GPFM compared to other model. Nevertheless, we also recognize that current off-the-shelf foundation models still exhibit potential in specific tasks, such as Phikon for report generation using TCGA data. This illustrates that the proposed unified knowledge distillation approach is not perfect and has room for improvement. Future research should concentrate on developing sophisticated methodologies to effectively distill and incorporate expert knowledge into one model, maximizing their potential across a broader spectrum of tasks. For example, further expanding the model’s parameter size to enhance its adaptability, facilitating a more comprehensive assimilation of knowledge from diverse foundation models. Additionally, the current GPFM is an unimodal foundation model, which limits its ability to effectively handle cross-modal tasks such as VQA. Given the prevalence of multimodal data in pathology, encompassing WSIs, reports, and genomic data, the development of a multi-modal pathology foundation model is more attractive. Such a model would be more adept at integrating heterogeneous information, offering a more holistic understanding of patient data and enhancing diagnostic accuracy. Last, the GPFM and existing foundation models function as ROI-level feature extractors. However, the majority of CPath tasks demand analysis at the whole-slide level. Recent advancements in long sequence modeling techniques [67, 68], exemplified by MambaMIL [69], have demonstrated success in WSI classification tasks. Developing whole slide-level foundation model could avoid the reliance on MIL frameworks, potentially leading to significant strides in CPath by providing more contextually rich interpretations of tissues.
@@ 4.1 Foundation Model Pretraining
CPath has emerged as a groundbreaking field that synergizes the power of AI with the expertise of pathologists, revolutionizing the practice of diagnosing and analyzing diseases. At the core of this transformative discipline lies the foundation model, which serves as the backbone for a wide range of applications in pathology. While there exist some readily available foundation models such as Ctranspath (pretrained on 32K TCGA slides) [37] and UNI [33] (pretrained on 100K private slides), the utilization of public data remains incomplete, and the evaluation of these models in CPath tasks is inadequate. The limited diversity of primary sites in the pretraining slides also restricts the adaptability of current foundation models for public CPath benchmarks. To facilitate the advancement of CPath, we meticulously curated a comprehensive dataset comprising 47 publicly available histopathology datasets, encompassing a wide spectrum of 34 distinct tissue types for pretraining and downstream tasks evaluation. The primary sites of tissues are reported in Extended Data Table 41. Leveraging this large-scale dataset, we developed a self-supervised learning approach with unified knowledge distillation to construct a foundation model that surpasses existing models.
@@
Dataset Preparation. To boost the performance of foundation models, the diverse datasets with various tissues are necessary. We have collected over 33 datasets as depicted in Extended Data Table 43 (from row 1 to row 33). To process WSIs, we employed the OpenSlide [70] and CLAM toolkit [71] to find all non-overlapping 512$\times$512 patches at level 0 that contain tissues. It is worth noting that we did not scale the patches to a uniform resolution, opting instead to use the original resolution of each WSI. This approach was implemented to increase the robustness of the foundation models to varying resolutions. For datasets that only contain ROI images, we extracted non-overlapping 512$\times$512 patches as well. Upon processing all 33 datasets, we obtained a comprehensive dataset, as presented in Extended Data Table 40. The dataset used for pretraining consists of 72,280 WSIs and a total of 190,212,668 patches.
@@
Pretraining with Self and Expert Knowledge Distillation. In CPath, current foundation models typically rely on state-of-the-art self-supervised pretraining (SSL) methods, such as DINOv2 [56] and iBOT [72]. These methods are applied directly to either private or public datasets. For instance, Phikon [32] is constructed based on 6,093 TCGA slides using iBOT, while UNI is built upon approximately 100,000 private and public slides using DINOv2. Due to its larger training dataset and more powerful SSL methods, UNI outperforms Phikon in various tasks. However, UNI still lags behind other foundation models in tasks related to text analysis and survival analysis due to its pretraining strategy and limited coverage of primary sites. To address the limitations of current foundation models and further enhance their performance, we propose a novel pretraining strategy involving Unified Knowledge Distillation. The framework of our proposed pretraining method is illustrated in Figure 1.c. Similar to DINOv2, we employ teacher-student networks with masking image modeling (MIM) loss [73] and DINO (self-distillation) [74, 72] loss to optimize the student network. Specifically, given an input image $x$, we obtain two augmented views, $u$ and $v$. Random masking is then applied to both $u$ and $v$, resulting in masked views, $\hat{u}$ and $\hat{v}$. For the MIM objective, the student network takes $\hat{u}$ and $\hat{v}$ as inputs and aims to predict the masked tokens. With the DINO objective, we first crop $n$ additional local views, $w_{i}$, and extract encoded class ([CLS]) tokens using the student network. Next, we obtain the [CLS] tokens of the global views ($u$ and $v$) using the teacher network. Finally, we compute the cross-entropy loss between the local views and global views’ [CLS] tokens. However, this strategy fails to leverage the knowledge from existing vision foundation models, such as UNI and vision-language foundation models like CONCH [35], which restricts their applicability to different tissue types. To transfer the knowledge from established pathology foundation models, we introduce a Expert Knowledge Distillation module to distill knowledge into the student network [38, 75]. Specifically, we adopt three foundation models that excel in classification tasks (UNI), survival analysis tasks (Phikon), and visual question answering tasks (CONCH). To achieve distillation, we use the student network to encode the global views $u$ and $v$ and extract the [CLS] and [PATCH] tokens. Additionally, we employ the off-the-shelf foundation models (UNI, Phikon, CONCH) to obtain their respective [CLS] and [PATCH] tokens. For aligning the class tokens, we utilize cosine similarity. As for the patch token alignment, we employ both cosine similarity and smooth L1 distance. The pseudo-code for this process is outlined in Algorithm 1. The hyperparameters used in the pretraining phase are provided in Extended Data Table 38. Once the student network is updated, we adopt Exponential Moving Average (EMA) to update the teacher network (GPFM).
@@ 4.2 WSI Classification
To evaluate the performance of the MIL model, we assess the balanced accuracy, weighted F1 score, and AUC, which consider the class imbalance present in the dataset. Our experiments encompass 12 pathology WSI classification tasks, including: (1) breast cancer metastasis detection, (2) coarse-grained breast carcinoma subtyping, (3) fine-grained breast carcinoma subtyping, (4) lobular & ductal carcinoma subtyping, (5) ovarian cancer subtyping, (6) renal cell carcinoma (RCC) subtyping based on TCGA, (7) non-small cell lung cancer subtyping based on TCGA, (8) non-small cell lung cancer subtyping based on CPTAC, (9) prostate cancer grade assessment, (10) TP53 mutation prediction on LUAD (TCGA), (11) Brain tumor subtyping, and (12) glioma IDH1 mutation prediction. The results of our experiments are presented in Extended Data Tables 3 to 14. In Table 2, we showcase the average performance across all 12 WSI classification tasks, demonstrating that our method achieves the best performance.
@@
CPTAC-NSCLC for NSCLC Subtyping (2 classes). This dataset was adopted for the subtyping of NSCLC using lung squamous cell carcinoma (LSCC) and LUAD WSIs sourced from CPTAC data portal[79]. After preprocessing, four LSCC slides and one LUAD slide are excluded due to a limited number of tissues, resulting in a total of 1,077 LSCC slides and 1,136 LUAD slides available for further analysis. For training and evaluation, the train-validation-test cohort is label-stratified in a ratio of 7:1:2, resulting in 1549 slides for training, 222 slides for validation, and 442 slides for testing. The experimental results are reported in Extended Data Table 4.
@@ 4.4 ROI Classification
For patch-level tissue classification tasks, we evaluate the transfer performance and representation ability of different foundation models using a linear probe, inspired by the approach employed in DINOv2. [56, 84]. Initially, we extract features from the images using the pretrained foundation models. Subsequently, we employ a linear layer for performing classification. To optimize the model, we utilize AdamW [85] with an initial learning rate of 5e-4 and weight decay of 1e-5. Additionally, we incorporate a cosine annealing scheduler to update the learning rate during training [86]. In order to obtain the best model, we set the maximum number of epochs to 3000 and implement early stopping with a patience of 100 epochs. For ensuring fair comparison, we maintain a consistent batch size of 256 across all methods.
To evaluate the performance of patch-level tissue classification, we consider the impact of class imbalance in the dataset and assess the metrics of balanced accuracy, weighted F1 score, and AUC. These metrics provide comprehensive insights into the classification performance, accounting for both accuracy and the ability to handle imbalanced class distributions. Specifically, we compare the foundation models across 12 tasks, including: (1) Colorectal cancer (CRC) tissue classification, (2) Clear cell colorectal cancer (CCRCC) tissue classification, (3) Breast cancer tissue classification, (4) CRC polyp classification, (5) Microsatellite instability screening, (6) PanCancer tissue classification, (7) Tumor infiltrating lymphocyte classification, (8) Esophageal carcinoma subtyping, (9) Metastatic tissue classification, (10) Lung adenocarcinoma tissue classification, (11) Breast cancer image classification, and (12) Colon tissue classification. For all experiments in this section, we estimate the model performance using non-parametric bootstrapping with 1,000 bootstrap replicates. We employ Torchmetrics [87] for bootstrapping sampling and obtain the mean and standard deviation of the metrics. The experimental results are presented in Extended Data Table 21 to Extended Data Table 32. Furthermore, we report the average performance of the patch-level tissue classification results arcoss 12 tasks in Table 20, demonstrating the superior performance of GPFM.
@@
CCRCC-TCGA-HEL for CCRCC Tissue Classification (4 classes). This dataset [88] comprises a total of 52,713 region of interest (ROI) images, each with dimensions of 300$\times$300 pixels. The dataset encompasses six distinct categories, namely: renal cancer (cancer, 13,057 ROIs), normal renal tissue (normal, 8,652 ROIs), stromal tissue (stroma, 5,460 ROIs), red blood cells (blood, 996 ROIs), empty background (empty, 16,026 ROIs), and other textures, including necrotic, torn, and adipose tissue (other, 8,522 ROIs). The image tiles were selected at random from two sources: the TCGA-KIRC WSIs and the Helsinki datasets. For training and evaluation, we focused on four specific categories: cancer, stroma, normal, and blood. This decision was made due to the potential ambiguities associated with the "$other"$ category and the lack of meaningful information conveyed by the "$empty"$ category. We randomly shuffle the samples and set the train-test split as 22530:5635 ratio. The experimental results are shown in Extended Data Table 22.
@@
ESCA for Esophageal Carcinoma Subtyping (11 classes). This dataset [95] comprises 367,229 images with size of 256$\times$256 pixels. These patches were obtained from 320 H&E WSIs of esophageal adenocarcinoma and adenocarcinoma of the esophagogastric junction, specifically, 22 slides from University Hospital Cologne (UKK), 62 slides from Landesklinikum Wiener Neustadt (WNS), 22 slides from TCGA, and 214 slides from the University Hospital Berlin Charite (CHA). These images were annotated and labeled with one of eleven classes: adventitia (71,131 ROIs), lamina propria mucosae (2,173 ROIs), muscularis mucosae (2,951 ROIs), muscularis propria (83,358 ROIs), regression tissue (56,490 ROIs), mucosa gastric (44,416 ROIs), muscosa oesophagus (18,561 ROIs), submucosa (22,117 ROIs), submucosal glands (1,516 ROIs), tumor (63,863 ROIs), and ulceration (753 ROIs). For training and evaluation, we adopt CHA dataset, containing 178,187 ROIs, as the training set, and we combined the UKK, WNS, and TCGA datasets as single testing cohort consisting of 189,142 ROIs. In our experiment, all images were resized to 224 × 224 pixels to ensure consistency, the experimental result is shown in Extended Data Table 28.
@@
WSSS4LUAD for Lung Adenocarcinoma Tissue Classification (3 classes). This dataset [97, 98] was collected from Guangdong Provincial People’s Hospital (GDPH) and TCGA. It consists of 10,091 images with following three common and meaningful tissue types: tumor epithelial tissue (6,579 ROIs), tumor-associated stroma tissue (1,680 ROIs), and normal tissue (1,832 ROIs). It is worth noting that, in WSSS4LUAD dataset, one image may belong to several categories. To avoid ambiguity, we only choose one label for each image based on the order of diagnosability (i.e., from tumor epithelial tissue to normal tissue). For training and evaluation, all images were resized to 224$\times$224 pixels and we label-stratified the train-test with a ratio of 8:2 (8,072:2019 ROIs). The experimental results are presented in Extended Data Table 30.
@@
In the linear probe evaluation tasks, we extract semantically-rich features using different foundation models and then construct task-specific classifier. These features are not only applicable for supervised learning but also prove to be valuable for image-to-image retrieval. The primary goal of this application is to retrieve images that share the same class label as a given query image, thereby facilitating efficient image retrieval. To initiate the pathological tissue image retrieval process, we begin by embedding all images using pretrained foundation models. Next, each image in the test set is treated as a query and compared against the images in the training set. To ensure that all features have a comparable impact on the computation of similarity, we independently normalize each feature component to the range [0, 1] [101]. This normalization process involves calculating the mean and variance of the training set features, which are then used to normalize both the training and testing features.

### v2

@@ 3 Discussion
This study also has some limitations. While we established a most comprehensive benchmark covering 39 tasks and demonstrated that the GPFM was the most generalizable foundation model, the external validation was limited to only seven tasks. This set of external validation tasks could be expanded in future work to further assess the generalizability of the models. It is worth highlighting that, across the seven external validation tasks, the GPFM achieved an average rank of 1.57, while the second-best performing model, UNI, achieved an average rank of 2.79. This result is consistent with the conclusions drawn from the experiments across the entire benchmark, highlighting the superior generalizability of GPFM compared to other models. Nevertheless, we also recognize that current off-the-shelf foundation models still exhibit potential in specific tasks, such as Phikon for report generation using TCGA data. This illustrates that the proposed unified knowledge distillation approach is not perfect and has room for improvement. Future research should concentrate on developing sophisticated methodologies to effectively distill and incorporate expert knowledge into one model, maximizing their potential across a broader spectrum of tasks. For example, further expanding the model’s parameter size to enhance its adaptability, facilitating a more comprehensive assimilation of knowledge from diverse foundation models. Additionally, the current GPFM is an unimodal foundation model, which limits its ability to effectively handle cross-modal tasks such as VQA. Given the prevalence of multimodal data in pathology, encompassing WSIs, reports, and genomic data, the development of a multi-modal pathology foundation model is more attractive. Such a model would be more adept at integrating heterogeneous information, offering a more holistic understanding of patient data and enhancing diagnostic accuracy. Last, the GPFM and existing foundation models function as ROI-level feature extractors. However, the majority of CPath tasks demand analysis at the whole-slide level. Recent advancements in long sequence modeling techniques [67, 68], exemplified by MambaMIL [69], have demonstrated success in WSI classification tasks. Developing whole slide-level foundation model could avoid the reliance on MIL frameworks, potentially leading to significant strides in CPath by providing more contextually rich interpretations of tissues.
@@ 4.1 Foundation Model Pretraining
CPath has emerged as a groundbreaking field that synergizes the power of AI with the expertise of pathologists, revolutionizing the practice of diagnosing and analyzing diseases. At the core of this transformative discipline lies the foundation model, which serves as the backbone for a wide range of applications in pathology. While there exist some readily available foundation models such as Ctranspath (pretrained on 32K TCGA slides) [37] and UNI [33] (pretrained on 100K private slides), the utilization of public data remains incomplete, and the evaluation of these models in CPath tasks is inadequate. The limited diversity of primary sites in the pretraining slides also restricts the adaptability of current foundation models for public CPath benchmarks. To facilitate the advancement of CPath, we meticulously curated a comprehensive dataset comprising 47 publicly available histopathology datasets, encompassing a wide spectrum of 34 distinct tissue types for pretraining and downstream task evaluation. The primary sites of tissues are reported in Extended Data Table 41. Leveraging this large-scale dataset, we developed a self-supervised learning approach with unified knowledge distillation to construct a foundation model that surpasses existing models.
@@
Dataset Preparation. To boost the performance of foundation models, diverse datasets with various tissues are necessary. We have collected over 33 datasets as depicted in Extended Data Table 43 (from row 1 to row 33). To process WSIs, we employed the OpenSlide [70] and CLAM toolkit [71] to find all non-overlapping 512$\times$512 patches at level 0 that contain tissues. It is worth noting that we did not scale the patches to a uniform resolution, opting instead to use the original resolution of each WSI. This approach was implemented to increase the robustness of the foundation models to varying resolutions. For datasets that only contain ROI images, we extracted non-overlapping 512$\times$512 patches as well. Upon processing all 33 datasets, we obtained a comprehensive dataset, as presented in Extended Data Table 40. The dataset used for pretraining consists of 72,280 WSIs and a total of 190,212,668 patches.
@@
Pretraining with Self and Expert Knowledge Distillation. In CPath, current foundation models typically rely on state-of-the-art self-supervised pretraining (SSL) methods, such as DINOv2 [56] and iBOT [72]. These methods are applied directly to either private or public datasets. For instance, Phikon [32] is constructed based on 6,093 TCGA slides using iBOT, while UNI is built upon approximately 100,000 private and public slides using DINOv2. Due to its larger training dataset and more powerful SSL methods, UNI outperforms Phikon in various tasks. However, UNI still lags behind other foundation models in tasks related to text analysis and survival analysis due to its pretraining strategy and limited coverage of primary sites. To address the limitations of current foundation models and further enhance their performance, we propose a novel pretraining strategy involving Unified Knowledge Distillation. The framework of our proposed pretraining method is illustrated in Figure 1.c. Similar to DINOv2, we employ teacher-student networks with masking image modeling (MIM) loss [73] and DINO (self-distillation) [74, 72] loss to optimize the student network. Specifically, given an input image $x$, we obtain two augmented views, $u$ and $v$. Random masking is then applied to both $u$ and $v$, resulting in masked views, $\hat{u}$ and $\hat{v}$. For the MIM objective, the student network takes $\hat{u}$ and $\hat{v}$ as inputs and aims to predict the masked tokens. With the DINO objective, we first crop $n$ additional local views, $w_{i}$, and extract encoded class ([CLS]) tokens using the student network. Next, we obtain the [CLS] tokens of the global views ($u$ and $v$) using the teacher network. Finally, we compute the cross-entropy loss between the local views and global views’ [CLS] tokens. However, this strategy fails to leverage the knowledge from existing vision foundation models, such as UNI and vision-language foundation models like CONCH [35], which restricts their applicability to different tissue types. To transfer the knowledge from established pathology foundation models, we introduce an Expert Knowledge Distillation module to distill knowledge into the student network [38, 75]. Specifically, we adopt three foundation models that excel in classification tasks (UNI), survival analysis tasks (Phikon), and visual question answering tasks (CONCH). To achieve distillation, we use the student network to encode the global views $u$ and $v$ and extract the [CLS] and [PATCH] tokens. Additionally, we employ the off-the-shelf foundation models (UNI, Phikon, CONCH) to obtain their respective [CLS] and [PATCH] tokens. For aligning the class tokens, we utilize cosine similarity. As for the patch token alignment, we employ both cosine similarity and smooth L1 distance. The pseudo-code for this process is outlined in Algorithm 1. The hyperparameters used in the pretraining phase are provided in Extended Data Table 38. Once the student network is updated, we adopt the Exponential Moving Average (EMA) to update the teacher network (GPFM).
@@ 4.2 WSI Classification
To evaluate the performance of the MIL model, we assess the balanced accuracy, weighted F1 score, and AUC, which consider the class imbalance present in the dataset. Our experiments encompass 12 pathology WSI classification tasks, including (1) breast cancer metastasis detection, (2) coarse-grained breast carcinoma subtyping, (3) fine-grained breast carcinoma subtyping, (4) lobular & ductal carcinoma subtyping, (5) ovarian cancer subtyping, (6) renal cell carcinoma (RCC) subtyping based on TCGA, (7) non-small cell lung cancer subtyping based on TCGA, (8) non-small cell lung cancer subtyping based on CPTAC, (9) prostate cancer grade assessment, (10) TP53 mutation prediction on LUAD (TCGA), (11) Brain tumor subtyping, and (12) glioma IDH1 mutation prediction. The results of our experiments are presented in Extended Data Tables 3 to 14. In Table 2, we showcase the average performance across all 12 WSI classification tasks, demonstrating that our method achieves the best performance.
@@
CPTAC-NSCLC for NSCLC Subtyping (2 classes). This dataset was adopted for the subtyping of NSCLC using lung squamous cell carcinoma (LSCC) and LUAD WSIs sourced from the CPTAC data portal[79]. After preprocessing, four LSCC slides and one LUAD slide are excluded due to a limited number of tissues, resulting in a total of 1,077 LSCC slides and 1,136 LUAD slides available for further analysis. For training and evaluation, the train-validation-test cohort is label-stratified in a ratio of 7:1:2, resulting in 1549 slides for training, 222 slides for validation, and 442 slides for testing. The experimental results are reported in Extended Data Table 4.
@@ 4.4 ROI Classification
For patch-level tissue classification tasks, we evaluate the transfer performance and representation ability of different foundation models using a linear probe, inspired by the approach employed in DINOv2. [56, 84]. Initially, we extract features from the images using the pretrained foundation models. Subsequently, we employ a linear layer for performing classification. To optimize the model, we utilize AdamW [85] with an initial learning rate of 5e-4 and weight decay of 1e-5. Additionally, we incorporate a cosine annealing scheduler to update the learning rate during training [86]. In order to obtain the best model, we set the maximum number of epochs to 3000 and implemented early stopping with patience of 100 epochs. For ensuring fair comparison, we maintain a consistent batch size of 256 across all methods.
To evaluate the performance of patch-level tissue classification, we consider the impact of class imbalance in the dataset and assess the metrics of balanced accuracy, weighted F1 score, and AUC. These metrics provide comprehensive insights into the classification performance, accounting for both accuracy and the ability to handle imbalanced class distributions. Specifically, we compare the foundation models across 12 tasks, including (1) Colorectal cancer (CRC) tissue classification, (2) Clear cell colorectal cancer (CCRCC) tissue classification, (3) Breast cancer tissue classification, (4) CRC polyp classification, (5) Microsatellite instability screening, (6) PanCancer tissue classification, (7) Tumor infiltrating lymphocyte classification, (8) Esophageal carcinoma subtyping, (9) Metastatic tissue classification, (10) Lung adenocarcinoma tissue classification, (11) Breast cancer image classification, and (12) Colon tissue classification. For all experiments in this section, we estimate the model performance using non-parametric bootstrapping with 1,000 bootstrap replicates. We employ Torchmetrics [87] for bootstrapping sampling and obtain the mean and standard deviation of the metrics. The experimental results are presented in Extended Data Table 21 to Extended Data Table 32. Furthermore, we report the average performance of the patch-level tissue classification results across 12 tasks in Table 20, demonstrating the superior performance of GPFM.
@@
CCRCC-TCGA-HEL for CCRCC Tissue Classification (4 classes). This dataset [88] comprises a total of 52,713 regions of interest (ROI) images, each with dimensions of 300$\times$300 pixels. The dataset encompasses six distinct categories, namely: renal cancer (cancer, 13,057 ROIs), normal renal tissue (normal, 8,652 ROIs), stromal tissue (stroma, 5,460 ROIs), red blood cells (blood, 996 ROIs), empty background (empty, 16,026 ROIs), and other textures, including necrotic, torn, and adipose tissue (other, 8,522 ROIs). The image tiles were selected at random from two sources: the TCGA-KIRC WSIs and the Helsinki datasets. For training and evaluation, we focused on four specific categories: cancer, stroma, normal, and blood. This decision was made due to the potential ambiguities associated with the "$other"$ category and the lack of meaningful information conveyed by the "$empty"$ category. We randomly shuffle the samples and set the train-test split as a 22530:5635 ratio. The experimental results are shown in Extended Data Table 22.
@@
ESCA for Esophageal Carcinoma Subtyping (11 classes). This dataset [95] comprises 367,229 images with size of 256$\times$256 pixels. These patches were obtained from 320 H&E WSIs of esophageal adenocarcinoma and adenocarcinoma of the esophagogastric junction, specifically, 22 slides from University Hospital Cologne (UKK), 62 slides from Landesklinikum Wiener Neustadt (WNS), 22 slides from TCGA, and 214 slides from the University Hospital Berlin Charite (CHA). These images were annotated and labeled with one of eleven classes: adventitia (71,131 ROIs), lamina propria mucosae (2,173 ROIs), muscularis mucosae (2,951 ROIs), muscularis propria (83,358 ROIs), regression tissue (56,490 ROIs), mucosa gastric (44,416 ROIs), muscosa oesophagus (18,561 ROIs), submucosa (22,117 ROIs), submucosal glands (1,516 ROIs), tumor (63,863 ROIs), and ulceration (753 ROIs). For training and evaluation, we adopted CHA dataset, containing 178,187 ROIs, as the training set, and we combined the UKK, WNS, and TCGA datasets as a single testing cohort consisting of 189,142 ROIs. In our experiment, all images were resized to 224 × 224 pixels to ensure consistency, the experimental result is shown in Extended Data Table 28.
@@
WSSS4LUAD for Lung Adenocarcinoma Tissue Classification (3 classes). This dataset [97, 98] was collected from Guangdong Provincial People’s Hospital (GDPH) and TCGA. It consists of 10,091 images with the following three common and meaningful tissue types: tumor epithelial tissue (6,579 ROIs), tumor-associated stroma tissue (1,680 ROIs), and normal tissue (1,832 ROIs). It is worth noting that, in WSSS4LUAD dataset, one image may belong to several categories. To avoid ambiguity, we only choose one label for each image based on the order of diagnosability (i.e., from tumor epithelial tissue to normal tissue). For training and evaluation, all images were resized to 224$\times$224 pixels and we label-stratified the train-test with a ratio of 8:2 (8,072:2019 ROIs). The experimental results are presented in Extended Data Table 30.
@@
In the linear probe evaluation tasks, we extract semantically-rich features using different foundation models and then construct a task-specific classifier. These features are not only applicable for supervised learning but also prove to be valuable for image-to-image retrieval. The primary goal of this application is to retrieve images that share the same class label as a given query image, thereby facilitating efficient image retrieval. To initiate the pathological tissue image retrieval process, we begin by embedding all images using pretrained foundation models. Next, each image in the test set is treated as a query and compared against the images in the training set. To ensure that all features have a comparable impact on the computation of similarity, we independently normalize each feature component to the range [0, 1] [101]. This normalization process involves calculating the mean and variance of the training set features, which are then used to normalize both the training and testing features.